# IRL-VLA: Training an Vision-Language-Action Policy via Reward World Model

### v2

@@ 3.4 Reinforcement Learning with RWM
Following DiffusionDrive [22], the diffusion policy $\pi_{\theta}$ can be viewed as an internal Markov decision process that starts from Gaussian noise and gradually denoises to produce an action sequence. Concretely, we sample $G$ trajectories and obtain their diffusion chains. The diffusion chain of a single trajectory is represented as
$\small\mathbf{x}=\bigl{(}x_{T},\,x_{T-1},\,\dots,\,x_{0}\bigr{)},$ (1)
where $T$ is the total number of denoising steps. For this chain:
$\small x_{T}\sim\mathcal{N}(0,\mathbf{I}),x_{t-1}\sim\pi_{\theta}\bigl{(}x_{t-1}\mid x_{t}\bigr{)},\quad t=T,T-1,\dots,1.$ (2)
Trajectories reward $R_{epdms}$ are predicted by RMW, which evaluates each trajectory by NC, DAC, DDC, TLC, EP, TTC, LK, HC as sub-rewards and returns an EPDMS as reward $R_{epdms}$. We then compute the group-standardized advantage:
$\small\hat{A}_{i}\;=\;\frac{R_{epdms}-\mathrm{mean}\bigl{(}R_{1..G}\bigr{)}}{\sqrt{\mathrm{var}\bigl{(}R_{1..G}\bigr{)}}},\quad i=1,\dots,G.$ (3)
Each conditional step in the diffusion chain follows a Gaussian policy:
$\small\pi_{\theta}\bigl{(}x_{t-1}\mid x_{t}\bigr{)}=\mathcal{N}\!\Bigl{(}x_{t-1};\,\mu_{\theta}(x_{t},t),\,\sigma_{t}^{2}I\Bigr{)},$ (4)
where $\mu_{\theta}(x_{t},t)$ is the model-predicted mean and $\sigma_{t}^{2}I$ is the fixed covariance.
Thus, the probability density of the full chain $\mathbf{x}_{0:T}$ under $\pi_{\theta}$ is
$\small\log\pi_{\theta}\bigl{(}\mathbf{x}_{0:T}\bigr{)}=\sum_{t=1}^{T}\log\pi_{\theta}\bigl{(}x_{t-1}\mid x_{t}\bigr{)}.\qquad$ (5)
Finally, we compute the policy loss following [28, 24, 2], incorporating a behavior cloning loss to prevent collapse during exploration:
$\small L_{RL}\;=\;-\frac{1}{G}\sum_{i=1}^{G}\frac{1}{T}\sum_{t=1}^{T}\gamma^{\,t-1}\,\log\pi_{\theta}\bigl{(}x_{t-1}^{(i)}\mid x_{t}^{(i)}),\hat{A}_{i}$ (6)
where $\gamma$ is the discount coefficient (mitigating instability in early denoising steps), and $x_{t-1},x_{t}$ are sampled from the reference policy $\pi_{\mathrm{RL}}$.
Through RWM-assisted reinforcement learning, the diffusion planner learns to predict safe, comfortable trajectories via exploration, going beyond mere imitation to inject cognition into our framework.
The final policy optimization loss combines the reinforcement learning objective with a behavior cloning term to maintain stability and prevent catastrophic forgetting of the pretrained policy:

### v3

@@ 3.4 Reinforcement Learning with RWM
Inspired by diffusion-based planning methods [22, 19], our diffusion policy $\pi_{\theta}$ can be viewed as an intrinsic Markov decision process that starts from Gaussian noise and progressively denoises it to produce a sequence of actions. Following this paradigm, we generate a set of trajectories $T_{trajs}$ and record their complete diffusion processes. For a single trajectory, the diffusion progression is defined as
$\small\mathbf{T_{traj}}=\bigl{(}T_{traj\tau},T_{traj\tau-1},\dots,T_{traj0}\bigr{)},$ (1)
where $\tau$ denotes the total number of denoising steps.
In our framework, the RWM evaluates each trajectory using a multi-criteria scoring system, combining comfort metrics (EP, LK, HC) with safety metrics (NC, DAC, DDC, TLC, TTC). These are aggregated into an EPDMS-based score $r_{epdms}$. This score is combined with the value estimation from a critical network, and the advantage values are computed using GAE [24] as the $R_{epdms}=GAE(r_{epdms}+v_{epdms})$. Further, inspired by [25, 17], to improve stability across the batch, we adopt the group-standardized advantage the subsequent processing step involves derivation of group-standardized advantage values,
$\small\hat{A}_{i}\;=\;\frac{R_{epdms}-\mathrm{mean}\bigl{(}R_{1...T_{trajs}}\bigr{)}}{\sqrt{\mathrm{var}\bigl{(}R_{1..T_{trajs}}\bigr{)}}},\quad i=1,\dots,T_{trajs}.$ (2)
In diffusion progress, each conditional transition is modeled by a Gaussian policy [2]:
$\small\pi_{\theta}\left(x_{t-1}\mid x_{t},c\right)=\mathcal{N}\bigl{(}x_{t-1};,\mu_{\theta}(x_{t},c,t),\sigma_{t}^{2}I\bigr{)},$ (3)
where $\mu_{\theta}(x_{t},t,c)$ is the model-predicted mean and $\sigma_{t}^{2}I$ is the fixed covariance, $c$ include current vehicle states, semantic reasoning features, 3D reasoning features, and navigation command.
Inspired by recogdrive [17], The joint log-likelihood of the entire trajectory under this diffusion policy can therefore be expressed as
$\small\log\pi_{\theta}\bigl{(}\mathbf{x}_{0:T}\bigr{)}=\sum_{t=1}^{\tau}\log\pi_{\theta}\bigl{(}x_{t-1}\mid x_{t},c\bigr{)}.\qquad$ (4)
For policy optimization, we adopt the reinforcement learning via PPO algorithm. The reinforcement learning loss is formulated as follows:
$L_{RL}$ $=-\frac{1}{T_{trajs}}\sum_{i=1}^{T}\frac{1}{\tau}\sum_{t=1}^{\tau}\gamma^{\,t-1}\,\log\pi_{\theta}\bigl{(}x_{t-1}^{(i)}\mid x_{t}^{(i)},c\bigr{)}\,\hat{A}_{i}$
$\quad-D_{KL}\!\left[\pi_{\theta}\,\big{\|}\,\pi_{ref}\right]$ (5)
where $\gamma$ is the discount coefficient (mitigating instability in early denoising steps), and $x_{t-1},x_{t}$ are sampled from the reference policy $\pi_{\mathrm{RL}}$, and $D_{KL}\!\left[\pi_{\theta}\,\big{\|}\,\pi_{ref}\right]$ is the KL divergence between the reference policy $\pi_{ref}$ and current policy $\pi_{\theta}$.
By leveraging RWM-assisted reinforcement learning, the diffusion-based planner acquires the capability to generate safe and comfortable trajectories through active exploration, moving beyond simple imitation and introducing cognitive reasoning into the framework. The final policy optimization loss combines the reinforcement learning objective with a behavior cloning term to maintain stability and prevent catastrophic forgetting of the pretrained policy: